Zeta Corporation is considering a project that costs $18,000,000 and will increase the after-tax cash flows by $4,000,000 per year for five years. The company has total assets of $80,000,000 and total debt of $20,000,000. The borrowing rate is 10 percent, the corporate tax rate is 21%, the risk-free rate is 3%, the beta of the stock is 1.5, and the return on the market is 12 percent The percent of debt in the capital structure is The cost of equity is The weighted average cost of capital i:s The NPV of the project is Should Zeta undertake the project? Explain
Answer #1
1) The percent of debt in the capital structure is 20000000/80000000 = 25.00%
2) The cost of equity per CAPM = risk free rate+beta*(market return-risk free rate) = 3%+1.5*(12%-3%) = 16.50%
3) WACC = 10%*(1-21%)*25%+16.50%*75% = 14.35%
4) NPV = -18000000+4000000*(1.1435^5-1)/(0.1435*1.1435^5) = $   -43,82,406
5) Zeta should not undertake the project as the NPV is negative.